a bullet incident normally on a wooden plank , loses 1/10 th of its speed in passing through the plank . the least number of such planks required to stop the bullet is _____ .
 
a. 5
  1. 6
      1. 8

Grade:11

1 Answers

Vikas TU
7 years ago
Writing direct for n planks,
0^2 = v^2+ 2and
or
n = -v^2/2ad.........(1)
 
Now to pass through one plank,
the final velocity becomes,
v – v/10 = 9v/10 m/s.
therfore,
(9v/10)^2 = v^2 + 2ad................(2)
put value of 2ad in eqn. (1) and get the value of n.